OpenStudy (anonymous):

What is the equation of the line that passes through the point (1, 6) and has a slope of 3 y = 3x + 5 y = 3x + 3 y = 3x - 17 y = 3x - 5

OpenStudy (anonymous):

y = 3x + 5

OpenStudy (anonymous):

y-y1=m(x-x1) so plug in (1,6) in the equation y-6= 3(x-1) y-6=3x-3 y=3x-3+6 so y=3x+3

OpenStudy (anonymous):

the slope is m=3